OpenStudy (anonymous):

Alright, I'm calling it. Help me guys, you're my only hope. I swear, I will medal you till the cows come home if you'll only help me with this question. I will follow you throughout OpenStudy, and medal you on every question you answer. Just, for the love of Gallifrey, help me out. P.S. It's Trig.

OpenStudy (anonymous):

Using the value from the previous question, find the measure of angle Y and the length of side y. Hint: X was 89.9

OpenStudy (anonymous):

Wouldn't angle Y be 22.72?

OpenStudy (anonymous):

Please explain. Sorry, it's free response, so I need to understand. Although it makes more sense than my previous answers, which were either more or less than the other leg.

jimthompson5910 (jim_thompson5910):

Let the remaining angle be angle Z Angle Y is found by adding up the three angles and setting it equal to 180. Then you solve for Y X+Y+Z = 180 89.9+Y+67.38 = 180 Y = 180-(89.9+67.38) Y = 22.72 ------------------------------------------------------- Then use the law of sines to get sin(Z)/XY = sin(Y)/XZ sin(Z)/XY = sin(Y)/y sin(67.38)/24 = sin(22.72)/y now solve for y

OpenStudy (anonymous):

Well all I did was ad 89.9 and 67.38 because the degrees in a triangle always equal 180. when you get the answer 157.28, you just subtract that from 180 and get your missing answer

OpenStudy (anonymous):

Yeah....I got the angle fine...but finding y is harder I don't get it

OpenStudy (anonymous):

Sorry, @taylormaya , yeah, I thought you were talking about side length y. Yeah, I got the right answer for the angle..

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Thank the Lord Jesus in the high heavens, and the Queen Lolth in her palace in the deepest level of the Nine Hells - Glory be to God, and peace to his men on Earth.

OpenStudy (anonymous):

Thank you thank you thank you thank you thank you

OpenStudy (anonymous):

Hallelujah (fyi I'm not religious...yet....I just grew up with a Catholic grandmother, and a mom raised by said granmother..we pick up some mannerisms;)